# Quantum Rainbow Codes

### v2

@@ Abstract
We introduce rainbow codes, a novel class of quantum error correcting codes generalising colour codes and pin codes. Rainbow codes can be defined on any $D$-dimensional simplicial complex that admits a valid $(D+1)$-colouring of its 00-simplices. We study in detail the case where these simplicial complexes are derived from chain complexes obtained via the hypergraph product and, by reinterpreting these codes as collections of colour codes joined at domain walls, show that we can obtain code families with growing distance and number of encoded qubits as well as logical non-Clifford gates implemented by transversal application of $T$ and $T^{\dagger}$. By combining these techniques with the quasi-hyperbolic colour codes of Zhu et al. (arXiv:2310.16982) we obtain a family of codes with transversal non-Clifford gates and parameters $[\![n,O(n),O(log(n))]\!]$. This is the first example of a family of LDPC codes with linear rate, growing distance and transversal non-Clifford gates, which are necessary conditions for the magic-state distillation parameter $\gamma=\log_{d}(n/k)$ to be made arbitrarily small. In contrast to other recent constructions that achieve $\gamma\rightarrow 0$, our codes are natively defined on qubits, are LDPC, and have non-Clifford gates implementable by single-qubit (rather than entangling) physical operations, but are not asymptotically good.
@@ I Introduction
In this work we further generalise colour codes and pin codes to obtain a class of codes we call rainbow codes. These codes are obtained by identifying the low-weight logical operators of the pin codes and including some of them in the stabiliser group. We will see that there are a number of different choices for which operators to include (and which to remove) from the stabiliser group and these different choices produce codes with very different properties. We show that some of these choices, when combined with the hypergraph product tillich_quantum_2014 , can result in families of codes with growing distance and number of encoded qubits as well as transversal implementations of logical non-Clifford gates. We also show that these codes can be interpreted as joinings of Euclidean colour codes at domain walls between two copies of the topological phase. By combining this construction with the recently proposed quasi-hyperbolic codes of zhu2023non we can obtain codes with finite asymptotic rate, non-constant distance and transversal non-Clifford gates. The existence of such codes is a necessary condition for the magic-state yield parameter $\gamma=\log_{d}(n/k)$ bravyi_universal_2005 ; bravyi_magic-state_2012 to be made arbitrarily small, and until very recently the construction of such codes was an open problem, although their existence has now been demonstrated for the case of asymptotically good non-LDPC codes with gates implemented by transversal $CCZ$ wills_constant-overhead_2024 ; nguyen_good_2024 ; golowich_asymptotically_2024 . In contrast, our codes have transversal non-Clifford gates implemented by transversal $T/T^{\dagger}$, are LDPC, but are not asymptotically good. We note also that the logical action of our non-Clifford gate results in states with a complex entanglement structure and it is non-obvious how to use these states in magic state distillation procedures.
@@
We begin this paper by reviewing the constructions of quantum colour codes and pin codes then defining rainbow codes in Section II. We then consider rainbow codes defined via the hypergraph product in Section III and show how these codes can be interpreted as joinings of colour codes on manifolds. In Section IV we consider the action of transversal non-Clifford gates on these codes and identify the cases in which they can implement logical non-Clifford operations. Finally we present several examples of both finite-size and asymptotic rainbow codes, including the previously mentioned family with linear encoding rate, in Section V. We also describe various other properties and transformations of these codes in the appendices, such as a method for modifying rainbow codes to reduce physical qubit count and stabiliser weight while potentially preserving $k$ and $d$ (Appendix A), an unfolding map for some classes of rainbow code (Appendix B) and various algorithms enabling efficient construction of rainbow codes from chain complexes (Appendix C).
@@ II.1 Colour codes and pin codes
Definition 1.
Consider a length $D$ chain complex
$\mathcal{C}=C_{D}\overset{\delta_{D}}{\rightarrow}C_{D-1}\leavevmode\nobreak\ ...\leavevmode\nobreak\ C_{1}\overset{\delta_{1}}{\rightarrow}C_{0}$
which describes a cellulation of a manifold. The subcomplex
@@
We will also sometimes use the term “colour code lattice” to refer to the colourable graph $\mathcal{G}$ identified with $C_{(1,0)}$. In the case that $\mathcal{C}_{(1,0)}$ is a colour code lattice we can define a colour code bombin2006topological ; kubica2018abcs on $\mathcal{C}$ by choosing a pair of integers $x,z$ such that
@@
and then assigning a qubit to each 00-cell of $\mathcal{C}$ and an $X$/$Z$ stabiliser to each $x$-/$z$-cell. An example is shown in Fig. 1. Notice that the colouring of the $1$-cells also induces a colouring on the $D$-cells. The colourability and valency conditions of $\mathcal{G}$, as well as the fact the complex comes from a manifold, ensure that this forms a valid CSS code. When taking $x=D$ and $z=2$ colour codes admit transversal diagonal gates of the form $R_{D}=\mathrm{diag}(1,e^{i\pi/2^{D-1}})$, which belong to the $D$-level of the Clifford hierarchy bombin2007topological ; bombin2015gauge ; kubica2015universal .
@@
An alternative way to define a $D$-dimensional colour code is using a $D$-dimensional simplicial complex $\mathcal{K}$ that cellulates a manifold and admits a $(D+1)$-dimensional colouring of its 00-simplices (i.e. 00-simplices that are part of the same $1$-simplex must have different colours). To see that this is true, and that these two methods of defining colour codes are equivalent, notice that the cellulation described by $\mathcal{K}$ is dual 111recall that the dual of a cellulation replaces $D$-cells with 00-cells, $(D-1)$-cells with $1$-cells and so on. to a cellulation described by some $\mathcal{C}$ where $\mathcal{C}_{(1,0)}$ is a colour code lattice. This is because the valency condition of $\mathcal{G}$ means that 00-cells of $\mathcal{C}$ are mapped to $D$-simplices of $\mathcal{K}$, and the $(D+1)$-colouring of the $D$-cells of $\mathcal{C}$ translates to a $(D+1)$-colouring of the 00-simplices of $\mathcal{K}$. The colour code defined on $\mathcal{K}$ then has qubits on $D$-simplices and stabilisers on $(D-x)$- and $(D-z)$-simplices (in the sense that they are supported on all $D$-simplices that contain this $(D-x)$ or $(D-z)$ simplex). An example of this duality is also shown in Fig. 1.
@@
Figure 1: A square-octagon lattice can be used to define a colour code in two dimensions. Colours can be assigned to both edges and faces so that $c$-coloured faces are made from non-$c$-coloured edges. Qubits are assigned to vertices and stabilisers to faces. In the dual picture each qubit is assigned to a triangle ($2$-simplex) and stabilisers to vertices of these triangles (which inherit a colouring from the lattice).
@@
The definition via simplicial complexes is useful as a simplicial complex with $(D+1)$-colourable vertices can be easily obtained from any length $D$ chain complex via the following procedure. First, a 00-simplex is associated to every $i$-cell $c_{i}$ of the chain complex. Next a $1$-simplex is associated to every pair of cells $c_{i}$ and $c_{j}$ such that $c_{i}\in c_{j}$, where we define this inclusion to mean that $i<j$ and there exists some length $j-i$ sequence
$c_{i}\in\delta_{i+1}(c_{i+1}),\leavevmode\nobreak\ c_{i+1}\in\delta_{i+2}(c_{i+2}),\leavevmode\nobreak\ ...,\leavevmode\nobreak\ c_{j-1}\in\delta_{j}(c_{j}).$ (2)
@@
Intuitively this means that $c_{i}$ is a part of some higher dimensional cell $c_{j}$. For instance, if $c_{2}$ is a square then $c_{1}\in c_{2}$ means $c_{1}$ is an edge of this square while $c_{0}\in c_{2}$ means $c_{0}$ is a vertex of this square. Higher dimensional simplicies are similarly defined, so that an $m$-simplex is associated to a set of $m+1$ different cells $c_{i},c_{j},...,c_{k}$ where $c_{i}\in c_{j}\in...\in c_{k}$. $D$-simplices are then associated with sets containing exactly one cell of each dimension such that $c_{i}\in c_{i+1}$ for all $0\leq i<D$. These $D$-simplices are sometimes referred to as flags, and the 00-simplices of a flag are naturally $(D+1)$-coloured by the $(D+1)$ different dimensions of $i$-cell to which they correspond. In fact, the simplices shown in Fig. 1 are the flags of a square lattice where each red 00-simplex corresponds to a vertex of this lattice, each light blue 00-simplex corresponds to an edge and each dark blue 00-simplex to a square face.
@@
This method of constructing simplicial complexes was used in vuillot2022quantum to obtain generalised colour codes called pin codes. In the language of pin codes the set of all $D$-simplices containing a given $k$-simplex is called a $k$-pinned set, and stabilisers are associated to $(D-x)$- and $(D-z)$-pinned sets where $x$ and $z$ are subject to the same constraints as in Eq. 1. These codes correspond exactly to colour codes if the simplicial complex from which they are defined corresponds to a cellulation of a manifold, but in general this does not need to be the case. In particular, the hypergraph product of classical codes can be used to obtain a chain complex and then a simplicial complex via the method described previously. Pin codes defined on these complexes were observed to have very high encoding rates but only constant distance in most cases. In addition, the transversal non-Clifford gate of high-dimensional colour codes were not typically inherited by these pin codes prior to additional modifications.
@@ II.2 The simplex graph
Figure 2: a) A $2$-dimensional simplicial complex that does not correspond to a cellulation of a manifold and its associated simplex graph. Vertices of the simplex graph can be part of more than one edge of the same colour. b) A $2$-maximal and c), d) $2$-rainbow subgraphs of the simplex graph, along with the associated parts of the complex.
@@
Given a $D$-dimensional simplicial complex, $\mathcal{K}$, and a $(D+1)$-colouring of the 00-simplices of this complex, the corresponding simplex graph, $\mathcal{G}_{\mathcal{K}}$, is the graph such that
@@
  • there is an edge between vertices whenever the corresponding $D$-simplices share a $(D-1)$-simplex
  • each edge has a colour $c$, which is the unique colour of 00-simplex not contained in the $(D-1)$-simplex associated with this edge.
Note that to avoid confusion we will use the terms “vertex” and “edge” only when referring to objects of $\mathcal{G}_{\mathcal{K}}$. Vertices and edges of $\mathcal{K}$ will always be referred to as 00- and $1$-simplices respectively.
Perhaps the simplest example of a simplex graph is a 2D colour code lattice like Fig. 1, where we have a $2$-dimensional simplicial complex with a $3$-colouring defined on the 00-simplices. There is then one graph/lattice vertex for each $2$-simplex and edges join these vertices whenever the intersection of the corresponding $2$-simplices is a $1$-simplex. Each such intersection contains two colours of 00-simplex and the corresponding edge of the simplex graph is coloured with the third colour. More generally we have the following fact
@@
If a $D$-dimensional simplicial complex with $(D+1)$-colourable 00-simplices corresponds to a cellulation of a closed $D$-dimensional manifold then the associated simplex graph corresponds to a colour code lattice.
@@
Because the simplical complex cellulates a $D$-dimensional manifold a unique pair of $D$-simplices must meet at each $(D-1)$-simplex. The number of $(D-1)$-simplices contained in a $D$-simplex is $D+1$, so each vertex in the simplex graph is $(D+1)$-valent. A natural $(D+1)$-colouring of the edges of the graph is inherited from the colouring of the simplicial complex. These are exactly the requirements for a graph that can be identified with a colour code lattice. ∎
For simplicial complexes that do not correspond to cellulations of manifolds there is no limit on how many $D$-simplices can meet at a $(D-1)$-simplex and so vertices in the simplex graph are not generally $(D+1)$-valent and can be part of multiple edges of the same colour. An example of such a graph is shown in Fig. 2.
@@
An $S$-maximal subgraph, $\mathcal{M}^{k}$, of $\mathcal{G}_{\mathcal{K}}$ is a maximal connected subgraph containing only edges with colours in $S$
@@
We will also use the terms $k$-maximal and $k$-rainbow subgraph to refer to $S$-maximal or $S$-rainbow subgraphs for any choice of $S$ of size-$k$. Examples of both types of subgraph are shown in Fig. 2. Notice that $1$-maximal subgraphs must be cliques (fully connected subgraphs), as if multiple $D$-simplices meet at a common $(D-1)$-simplex their corresponding vertices in $\mathcal{G}_{\mathcal{K}}$ must all be connected by edges of the same colour.
These subgraphs can be used to define both colour codes and pin codes. For a colour code (i.e. a code defined from a simplicial complex that cellulates a manifold) $k$-maximal and $k$-rainbow subgraphs are equivalent since each vertex in the simplex graph is part of only one edge of each colour. Stabilisers of the colour code can then be thought of as being assigned to either $x$- and $z$-maximal subgraphs, $x$- and $z$-rainbow subgraphs or any mix of the two. For instance, every hexagonal face of the colour code lattice shown in Fig. 1 is both a $2$-maximal and $2$-rainbow subgraph. To see the connection to pin codes, notice that the set of vertices contained in a $k$-maximal subgraph is equivalent to the set of $D$-simplices contained in a $(D-k)$-pinned set. This is because vertices in the simplex graph connected by $c_{i}$-coloured edges correspond to $D$-simplices which differ only by a $c_{i}$-coloured 00-simplex. As a result, an $S$-maximal subgraph corresponds to a maximal set of $D$-simplices which differ from each other only by 00-simplices with colours in $S$, and so there is a common $(D-k)$-simplex shared by all $D$-simplices in this set. For example, all $2$-simplices associated with the $2$-maximal subgraph shown in Fig. 2 b) have a common $(2-2=0)$-simplex. We can therefore view pin codes as being defined by the assignment of $X$ and $Z$ stabilisers to all $x$- and $z$-maximal subgraphs of a simplex graph, and thus they are equivalent to colour codes when this simplex graph corresponds to a $D$-dimensional lattice.
@@
The (nontrivial) intersection of an $S_{1}$-maximal subgraph, $\mathcal{M}^{\lvert S_{1}\rvert}$, and an $S_{2}$-maximal subgraph, $\mathcal{M}^{\lvert S_{2}\rvert}$, is an $(S_{1}\cap S_{2})$-maximal subgraph (or multiple such subgraphs).
@@
By Lemma 3, for some $\mathcal{M}^{x}$ and $\mathcal{R}^{z}$ with nontrivial intersection we have $\mathcal{M}^{x}\cap\mathcal{R}^{z}=\mathcal{R}_{1}^{m}\cup\mathcal{R}_{2}^{m}\cup...$ for some $m$. Because $x+z\geq D+2$ (by Eq. 1) but there are only $D+1$ different colours in the graph the two subgraphs must have at least one colour of edge in common, and so $m\geq 1$. If $m>1$ we can choose some $T$ such that each $\mathcal{R}_{i}^{m}/T$ is a collection of disjoint $1$-rainbow subgraphs (which are just single edges connecting pairs of vertices), and otherwise $\mathcal{R}_{i}^{m}$ itself is a collection of disjoint $1$-rainbow subgraphs. $\mathcal{M}^{x}\cap\mathcal{R}^{z}$ must therefore contain an even number of vertices. ∎

### v3

@@ Abstract
We introduce rainbow codes, a novel class of quantum error correcting codes generalising colour codes and pin codes. Rainbow codes can be defined on any $D$-dimensional simplicial complex that admits a valid $(D+1)$-colouring of its 0-simplices. We study in detail the case where these simplicial complexes are derived from chain complexes obtained via the hypergraph product and, by reinterpreting these codes as collections of colour codes joined at domain walls, show that we can obtain code families with growing distance and number of encoded qubits as well as logical non-Clifford gates implemented by transversal application of $T$ and $T^{\dagger}$. By combining these techniques with the quasi-hyperbolic colour codes of Zhu et al. (arXiv:2310.16982) we obtain a family of codes with transversal non-Clifford gates and parameters $n,\Theta(n),\Theta(\log(n))]\!]$. This is the first example of a family of LDPC codes with linear rate, growing distance and transversal non-Clifford gates, which are necessary conditions for the magic-state distillation parameter $\gamma=\log_{d}(n/k)$ to be made arbitrarily small. In contrast to several other constructions that satisfy these requirements, our codes are natively defined on qubits, are LDPC, and have non-Clifford gates implementable by single-qubit (rather than entangling) physical operations, but are not asymptotically good.
@@ I Introduction
In this work we further generalise colour codes and pin codes to obtain a class of codes we call rainbow codes. These codes are obtained by identifying the low-weight logical operators of the pin codes and including some of them in the stabiliser group. We will see that there are a number of different choices for which operators to include (and which to remove) from the stabiliser group and these different choices produce codes with very different properties. We show that some of these choices, when combined with the hypergraph product tillich_quantum_2014, can result in families of codes with growing distance and number of encoded qubits as well as transversal implementations of logical non-Clifford gates. We also show that these codes can be interpreted as joinings of Euclidean colour codes at domain walls between two copies of the topological phase. By combining this construction with the recently proposed quasi-hyperbolic codes of zhu2023non we can obtain codes with finite asymptotic rate, non-constant distance and transversal non-Clifford gates. The existence of such codes is a necessary condition for the magic-state yield parameter $\gamma=\log_{d}(n/k)$ bravyi_universal_2005; bravyi_magic-state_2012 to be made arbitrarily small, and until very recently the construction of such codes was an open problem, although their existence has now been demonstrated for the case of asymptotically good non-LDPC codes with gates implemented by transversal $CCZ$ wills_constant-overhead_2024; nguyen_good_2024; golowich_asymptotically_2024. In contrast, our codes have transversal non-Clifford gates implemented by transversal $T/T^{\dagger}$, are LDPC, but are not asymptotically good. We note also that the logical action of our non-Clifford gate results in states with a complex entanglement structure and it is non-obvious how to use these states in magic state distillation procedures. Subsequent to the release of the first version of this work, a construction for LDPC codes that provably achieve $\gamma\rightarrow 0$ was identified by Golowich and Lin golowich2024quantumldpccodestransversal.
@@
We begin this paper by reviewing the constructions of quantum colour codes and pin codes then defining rainbow codes in Section˜II. We then consider rainbow codes defined via the hypergraph product in Section˜III and show how these codes can be interpreted as joinings of colour codes on manifolds. In Section˜IV we consider the action of transversal non-Clifford gates on these codes and identify the cases in which they can implement logical non-Clifford operations. Finally we present several examples of both finite-size and asymptotic rainbow codes, including the previously mentioned family with linear encoding rate, in Section˜V. We also describe various other properties and transformations of these codes in the appendices, such as a method for modifying rainbow codes to reduce physical qubit count and stabiliser weight while potentially preserving $k$ and $d$ (Appendix˜A), an unfolding map for some classes of rainbow code (Appendix˜B) and various algorithms enabling efficient construction of rainbow codes from chain complexes (Appendix˜C).
@@ II.1 Colour codes and pin codes
Definition 1 (Chain complex).
A length-$D$ chain complex is the data of $D+1$ vector spaces $C_{1},\ldots,C_{D}$ and linear maps $\delta_{1},\ldots,\delta_{D}$ with $\delta_{i}:C_{i}\rightarrow C_{i-1}$, called boundary maps, such that $\delta_{i}\circ\delta_{i+1}=0$ for all $i\in\{1,\ldots,D-1\}$. We write a chain complex with the following diagram:
$\mathcal{C}=C_{D}\overset{\delta_{D}}{\rightarrow}C_{D-1}\penalty 10000\ ...\penalty 10000\ C_{1}\overset{\delta_{1}}{\rightarrow}C_{0}$
The elements of each vector space $C_{k}$ are called $\bm{k}$-chains.
In the rest of this paper, we will only consider chain complexes defined with finite-dimensional vector spaces over the field $\mathbb{F}_{2}$. In this case, we can equip each vector space with a basis, and represent the boundary maps as binary matrices. We call each basis element of $C_{k}$ a $\bm{k}$-cell. A chain complex equipped with those bases can then be viewed as a $(D+1)$-partite graph, whose nodes are the $k$-cells, where $k$ determines the type of a node in the partition. The edges are then determined by the boundary maps $\delta_{k}$, seen as biadjacency matrices between nodes of type $k$ and k1k-1. Conversely, let $\mathcal{G}$ be a $(D+1)$-partite graph such that nodes of type $k$ are only connected to nodes of type k1k-1 and $k+1$, with a biadjacency matrix $\delta_{k}$ (for $1\leq k\leq D$). Then, if the biadjacency matrices satisfy $\delta_{k-1}\circ\delta_{k}=0$, we can associate a chain complex to the graph, by defining the vector spaces as $C_{k}=\mathbb{F}_{2}^{n_{k}}$ and the boundary maps as $\delta_{k}$, where $n_{k}$ is the number of nodes of type $k$ in $\mathcal{G}$.
A particular type of chain complex arises when considering the cellulation of a manifold. Given a smooth $D$-dimensional manifold $\mathcal{M}$, a cellulation is a covering of $\mathcal{M}$ by a set of $D$-dimensional polytopes $\{P_{1},\ldots,P_{n}\}$, sending each polytope to $\mathcal{M}$ using injective homeomorphisms $\phi_{i}:P_{i}\rightarrow\mathcal{M}$, such that for all $i,j$, $i\neq j$, $\phi(P_{i})$ and $\phi(P_{j})$ are either disjoint or overlap on exactly one of their $k$-faces for any $k$ breuckmann2018phdthesis. The $k$-faces of a $D$-dimensional polytope refer to its $k$-dimensional elements: vertices (0-faces), edges ($1$-faces), faces ($2$-faces), and so on. The image of a $k$-face of a polytope $P_{i}$ by $\phi_{i}$ is called a $k$-cell of the cellulation. To the cellulation of a manifold we can associate a chain complex, built from a multipartite graph called the Hasse diagram of the cellulation. It is defined as follows: each $k$-cell defines a node of type $k$, and there is an edge between a node of type $k$ and a node of type k1k-1 if and only if the associated $k$-cell contains the associated (k1)(k-1)-cell. The Hasse diagram is characterized by the following important lemma (breuckmann2018phdthesis, Lemma 2.9).
Lemma 1.
Consider a manifold cellulation and let $c_{i+1}$ and $c_{i-1}$ be $(i+1)$- and (i1)(i-1)-cells respectively. The number of $i$-cells connected to both $c_{i+1}$ and $c_{i-1}$ in the Hasse diagram is either zero or two.
Using this lemma, it is easy to prove that the Hasse diagram has the right properties to define a chain complex, namely that successive biadjacency matrices $\delta_{k}$ and $\delta_{k-1}$ satisfy $\delta_{k-1}\circ\delta_{k}=0$. Intuitively, this is equivalent to the statement that in a manifold cellulation, the boundary of the boundary of a cell is always zero. We call such a chain complex a cell complex.
Definition 2 (Cell complex).
A cell complex is a chain complex obtained from the Hasse diagram of the cellulation of a manifold. Each vector space $C_{k}$ of this complex is naturally equipped with a basis, corresponding to the $k$-cells of the cellulation.
Another type of chain complex we will use throughout this work is the simplicial complex.
Definition 3 (Simplicial complex).
A $D$-dimensional simplicial complex is a length-$D$ chain complex equipped with a basis of $k$-cells, also called $\bm{k}$-simplices, such that every $k$-simplex is connected to exactly $k$ (k1)(k-1)-simplices in the associated multipartite graph.
A particular kind of simplicial complex emerges from the cellulation of a manifold by simplices, that is, when all the polytopes $P_{i}$ are simplices. We will call such a complex a simplicial cell complex.
Equipped with those definitions, we are now ready to define colour codes and pin codes.
Definition 4 (Colour code lattice).
Consider a length-$D$ cell complex
$\mathcal{C}=C_{D}\overset{\delta_{D}}{\rightarrow}C_{D-1}\penalty 10000\ ...\penalty 10000\ C_{1}\overset{\delta_{1}}{\rightarrow}C_{0}$
The subcomplex
@@
We will use the term “colour code lattice” to refer interchangeably to the length-1 complex $C_{(1,0)}$ or the to the colourable graph $\mathcal{G}$ associated to it. In the case that $\mathcal{C}_{(1,0)}$ is a colour code lattice we can define a colour code bombin2006topological; kubica2018abcs on $\mathcal{C}$ by choosing a pair of integers $x,z$ such that
@@
and then assigning a qubit to each 0-cell of $\mathcal{C}$ and an XX/$Z$ stabiliser to each xx-/$z$-cell. An example is shown in Fig.˜1. Notice that the colouring of the $1$-cells also induces a colouring on the $D$-cells. The colourability and valency conditions of $\mathcal{G}$, as well as the fact the complex comes from a manifold, ensure that this forms a valid CSS code. When taking $x=D$ and $z=2$ colour codes admit transversal diagonal gates of the form $R_{D}=\mathrm{diag}(1,e^{i\pi/2^{D-1}})$, which belong to the $D$-level of the Clifford hierarchy bombin2007topological; bombin2015gauge; kubica2015universal.
@@
An alternative way to define a $D$-dimensional colour code is using a $D$-dimensional simplicial cell complex $\mathcal{K}$ that admits a $(D+1)$-dimensional colouring of its 0-simplices (i.e. 0-simplices that are part of the same $1$-simplex must have different colours). To see that this is true, and that these two methods of defining colour codes are equivalent, notice that the cellulation described by $\mathcal{K}$ is dual 111recall that the dual of a cellulation replaces $D$-cells with 0-cells, (D1)(D-1)-cells with $1$-cells and so on. to a cellulation described by some $\mathcal{C}$ where $\mathcal{C}_{(1,0)}$ is a colour code lattice. This is because the valency condition of $\mathcal{G}$ means that 0-cells of $\mathcal{C}$ are mapped to $D$-simplices of $\mathcal{K}$, and the $(D+1)$-colouring of the $D$-cells of $\mathcal{C}$ translates to a $(D+1)$-colouring of the 0-simplices of $\mathcal{K}$. The colour code defined on $\mathcal{K}$ then has qubits on $D$-simplices and stabilisers on (Dx)(D-x)- and (Dz)(D-z)-simplices (in the sense that they are supported on all $D$-simplices that contain this (Dx)(D-x) or (Dz)(D-z) simplex). An example of this duality is also shown in Fig.˜1.
@@
Figure 1: A square-octagon lattice can be used to define a colour code in two dimensions. Colours can be assigned to both edges and faces so that $c$-coloured faces are made from non-$c$-coloured edges. Qubits are assigned to vertices and stabilisers to faces. In the dual picture each qubit is assigned to a triangle (22-simplex) and stabilisers to vertices of these triangles (which inherit a colouring from the lattice).
@@
The definition via simplicial complexes is useful as a simplicial complex with $(D+1)$-colourable vertices can be easily obtained from any length $D$ chain complex via the following procedure. First, a 0-simplex is associated to every $i$-cell $c_{i}$ of the chain complex. Next a $1$-simplex is associated to every pair of cells $c_{i}$ and $c_{j}$ such that $c_{i}\in c_{j}$, where we define this inclusion to mean that $i<j$ and there exists some length jij-i sequence
$c_{i}\in\delta_{i+1}(c_{i+1}),\penalty 10000\ c_{i+1}\in\delta_{i+2}(c_{i+2}),\penalty 10000\ ...,\penalty 10000\ c_{j-1}\in\delta_{j}(c_{j}).$ (2)
@@
Intuitively this means that $c_{i}$ is a part of some higher dimensional cell $c_{j}$. For instance, if $c_{2}$ is a square then $c_{1}\in c_{2}$ means $c_{1}$ is an edge of this square while $c_{0}\in c_{2}$ means $c_{0}$ is a vertex of this square. Higher dimensional simplicies are similarly defined, so that an $m$-simplex is associated to a set of $m+1$ different cells $c_{i},c_{j},...,c_{k}$ where $c_{i}\in c_{j}\in...\in c_{k}$. $D$-simplices are then associated with sets containing exactly one cell of each dimension such that $c_{i}\in c_{i+1}$ for all $0\leq i<D$. These $D$-simplices are sometimes referred to as flags, and the 0-simplices of a flag are naturally $(D+1)$-coloured by the $(D+1)$ different dimensions of $i$-cell to which they correspond. In fact, the simplices shown in Fig.˜1 are the flags of a square lattice where each red 0-simplex corresponds to a vertex of this lattice, each light blue 0-simplex corresponds to an edge and each dark blue 0-simplex to a square face.
@@
This method of constructing simplicial complexes was used in vuillot2022quantum to obtain generalised colour codes called pin codes. In the language of pin codes the set of all $D$-simplices containing a given $k$-simplex is called a $k$-pinned set, and stabilisers are associated to (Dx)(D-x)- and (Dz)(D-z)-pinned sets where $x$ and $z$ are subject to the same constraints as in Eq.˜1. These codes correspond exactly to colour codes if the simplicial complex from which they are defined corresponds to a cellulation of a manifold, but in general this does not need to be the case. In particular, the hypergraph product of classical codes can be used to obtain a chain complex and then a simplicial complex via the method described previously. Pin codes defined on these complexes were observed to have very high encoding rates but only constant distance in most cases. In addition, the transversal non-Clifford gate of high-dimensional colour codes were not typically inherited by these pin codes prior to additional modifications.
@@ II.2 The simplex graph
Figure 2: a) A $2$-dimensional simplicial complex that does not correspond to a cellulation of a manifold and its associated simplex graph. Vertices of the simplex graph can be part of more than one edge of the same colour. b) A 22-maximal and c), d) 22-rainbow subgraphs of the simplex graph, along with the associated parts of the complex.
@@
Given a $D$-dimensional simplicial complex, $\mathcal{K}$, and a $(D+1)$-colouring of the 0-simplices of this complex, the corresponding simplex graph, $\mathcal{G}_{\mathcal{K}}$, is the graph such that
@@
  • there is an edge between vertices whenever the corresponding $D$-simplices share a (D1)(D-1)-simplex
  • each edge has a colour $c$, which is the unique colour of 0-simplex not contained in the (D1)(D-1)-simplex associated with this edge.
Note that to avoid confusion we will use the terms “vertex” and “edge” only when referring to objects of $\mathcal{G}_{\mathcal{K}}$. Vertices and edges of $\mathcal{K}$ will always be referred to as 0- and $1$-simplices respectively.
Perhaps the simplest example of a simplex graph is a 2D colour code lattice like Fig.˜1, where we have a $2$-dimensional simplicial complex with a $3$-colouring defined on the 0-simplices. There is then one graph/lattice vertex for each $2$-simplex and edges join these vertices whenever the intersection of the corresponding $2$-simplices is a $1$-simplex. Each such intersection contains two colours of 0-simplex and the corresponding edge of the simplex graph is coloured with the third colour. More generally we have the following fact
@@
Any $D$-dimensional simplicial cell complex with $(D+1)$-colourable 0-simplices has its associated simplex graph correspond to a colour code lattice.
@@
Because the simplicial cell complex cellulates a $D$-dimensional manifold, a unique pair of $D$-simplices must meet at each (D1)(D-1)-simplex. The number of (D1)(D-1)-simplices contained in a $D$-simplex is $D+1$, so each vertex in the simplex graph is $(D+1)$-valent. A natural $(D+1)$-colouring of the edges of the graph is inherited from the colouring of the simplicial complex. These are exactly the requirements for a graph that can be identified with a colour code lattice. ∎
For simplicial complexes that do not correspond to cellulations of manifolds there is no limit on how many $D$-simplices can meet at a (D1)(D-1)-simplex and so vertices in the simplex graph are not generally $(D+1)$-valent and can be part of multiple edges of the same colour. An example of such a graph is shown in Fig.˜2.
@@
An $S$-maximal subgraph, $\mathcal{M}^{k}$, of $\mathcal{G}_{\mathcal{K}}$ is a connected subgraph containing only edges with colours in $S$, and such that no more edges of colour in $S$ can be added to the subgraph while maintaining connectivity.
@@
We will also use the terms $k$-maximal and $k$-rainbow subgraph to refer to $S$-maximal or $S$-rainbow subgraphs for any choice of $S$ of size-$k$. Examples of both types of subgraph are shown in Fig.˜2. Notice that $1$-maximal subgraphs must be cliques (fully connected subgraphs), as if multiple $D$-simplices meet at a common (D1)(D-1)-simplex their corresponding vertices in $\mathcal{G}_{\mathcal{K}}$ must all be connected by edges of the same colour.
These subgraphs can be used to define both colour codes and pin codes. For a colour code (i.e. a code defined from a simplicial cell complex) $k$-maximal and $k$-rainbow subgraphs are equivalent since each vertex in the simplex graph is part of only one edge of each colour. Stabilisers of the colour code can then be thought of as being assigned to either $x$- and $z$-maximal subgraphs, $x$- and $z$-rainbow subgraphs or any mix of the two. For instance, every hexagonal face of the colour code lattice shown in Fig.˜1 is both a $2$-maximal and $2$-rainbow subgraph. To see the connection to pin codes, notice that the set of vertices contained in a $k$-maximal subgraph is equivalent to the set of $D$-simplices contained in a (Dk)(D-k)-pinned set. This is because vertices in the simplex graph connected by $c_{i}$-coloured edges correspond to $D$-simplices which differ only by a $c_{i}$-coloured 0-simplex. As a result, an $S$-maximal subgraph corresponds to a maximal set of $D$-simplices which differ from each other only by 0-simplices with colours in $S$, and so there is a common (Dk)(D-k)-simplex shared by all $D$-simplices in this set. For example, all $2$-simplices associated with the $2$-maximal subgraph shown in Fig.˜2 b) have a common (22=0)(2-2=0)-simplex. We can therefore view pin codes as being defined by the assignment of $X$ and $Z$ stabilisers to all $x$- and $z$-maximal subgraphs of a simplex graph, and thus they are equivalent to colour codes when this simplex graph corresponds to a $D$-dimensional lattice.
@@
The (non-empty) intersection of an $S_{1}$-maximal subgraph, $\mathcal{M}^{\lvert S_{1}\rvert}$, and an $S_{2}$-maximal subgraph, $\mathcal{M}^{\lvert S_{2}\rvert}$, is an $(S_{1}\cap S_{2})$-maximal subgraph (or multiple such subgraphs).
@@
By Lemma˜4, for some $\mathcal{M}^{x}$ and $\mathcal{R}^{z}$ with nontrivial intersection we have $\mathcal{M}^{x}\cap\mathcal{R}^{z}=\mathcal{R}_{1}^{m}\cup\mathcal{R}_{2}^{m}\cup$... for some $m$. Because $x+z\geq D+2$ (by Eq.˜1) but there are only $D+1$ different colours in the graph the two subgraphs must have at least one colour of edge in common, and so $m\geq 1$. If $m>1$ we can choose some $T$ such that each $\mathcal{R}_{i}^{m}/T$ is a collection of disjoint $1$-rainbow subgraphs (which are just single edges connecting pairs of vertices), and otherwise $\mathcal{R}_{i}^{m}$ itself is a collection of disjoint $1$-rainbow subgraphs. $\mathcal{M}^{x}\cap\mathcal{R}^{z}$ must therefore contain an even number of vertices. ∎